OpenStudy (rainbow_rocks03):

PLZ HELP ME I will medal+fan=you for BEST response. Find the total cost. purchase price = $2615; tax rate = 3% A. $78.45 B. $2615.03 C. $2618.00 D. $2693.45

OpenStudy (anonymous):

C (:

OpenStudy (calculusxy):

2615 x 0.03 = x 2615 + x = total cost

OpenStudy (rainbow_rocks03):

I added 78.45+2615=2693.45

OpenStudy (anonymous):

nooo

OpenStudy (calculusxy):

I think that's correct @rainbow_rocks03

OpenStudy (rainbow_rocks03):

2615*0.03=78.45

OpenStudy (rainbow_rocks03):

So I am correct @calculusxy ?

OpenStudy (anonymous):

no idiot

OpenStudy (calculusxy):

I think so.

OpenStudy (calculusxy):

Why not? @gage224

OpenStudy (rainbow_rocks03):

:'(

OpenStudy (majikdusty):

Your starting price is 2615. You need to find out how much 3% of that is. That is the amount you pay in tax. you must turn 3% into a decimal which is done simply by dividing by 100 OR just moving the decimal place over. 2615 * .03 = 78.45 Now you must add this to the starting total which is 2693.45 your answer is D

OpenStudy (calculusxy):

@MajikDUSTY Yeah that's what I thought too.

OpenStudy (rainbow_rocks03):

I knew it was D.

OpenStudy (rainbow_rocks03):

yay I got it right.

OpenStudy (calculusxy):

@rainbow_rocks03 You see... :)

OpenStudy (calculusxy):

Good job :)